# Uncovering flow and deformation regimes in the coupled fluid-solid vestibular system

### v1

@@ 7.1 Implications and applications
Here, it is worth considering the distinction between the soft cupula ($\kappa\ll 1$) and stiff cupula ($\kappa\gg 1$) regimes in terms of dimensional quantities. Recall that $\kappa$ is defined as $\kappa=B\mathcal{T}/(a^{2}R\mu)$, where $\mathcal{T}$ is the timescale for the head motion, $a$ and $R$ are respectively the small and large radii defining the canals, and $B$ is the cupula’s bending stiffness. We see that the soft cupula regime may be attained for fast movements (small $\mathcal{T}$), large canals ($a,R$ large) and of course soft cupulas in absolute terms (small $B$). The converse holds for the stiff cupula regime. Inserting parameter values for $a$, $R$, $B$ and $\mu$ into the transition value predicted by our model, $\kappa_{c}=16I_{4}\alpha(b)$, we obtain a critical value
$\mathcal{T}=\frac{a^{2}R\mu\kappa_{c}}{B}=\frac{12(1-\nu^{2})\,a^{2}R\mu\,\kappa_{c}}{E\,t_{h}^{3}}\sim\frac{a^{2}R\mu}{Et_{h}^{3}},$ (58)
which may be interpreted as a critical frequency of rotation below which the system responds to angular velocity, and above which the system responds to angular acceleration. Inserting typical values for a human adult, we compute a transition frequency of 0.27 Hz. Interestingly, human experiments with controlled oscillation frequencies have reported a maximum motion sickness when the frequency is around 0.2 Hz (Golding et al., 2001). Our analysis suggests an intriguing possible explanation for this maximal sickness at intermediate frequencies: intermediate frequencies correspond to motion for which the response of the cupular system follows neither the angular acceleration nor velocity. The “neural mismatch” hypothesis predicts that motion sickness is induced in situations where there is a disagreement between visual or vestibular cues and the information anticipated by the nervous system (Benson, 1990). Since the vestibular system may be expected to provide information that matches neither the true acceleration or the true velocity around the transition point we suggest that it may be the transition point between small and large $\kappa$ that causes motion sickness.
@@ 7.2 Limitations and future directions
Our model has been based on a number of approximations that have facilitated the analysis that we have presented here. Of these, perhaps the most important is our use of a thin plate model of the cupula — this is mathematically convenient but is likely to be an over-simplification of the true behaviour of the cupula. In particular, its validity would require the cupula to be slender; concrete measurements of the cupula’s aspect ratio are not available, meaning our thin-plate model of the cupula might not be appropriate in all cases. As an alternative strategy, we note that the strains are small, and so a solution of the Navier equations (2) that does not assume the thickness of the cupula to be small might be required. At the same time, we note that our existing analysis assumes that the cupula’s thickness is constant, but photographic evidence suggests this is likely not the case (Rabbitt et al., 2004).In particular, the cupula seems to be thinner in the centre and thicker towards the edge. Modelling this effect, via a plate equation with a spatially dependent thickness, might be enough to warrant further work with the plate model. Whilst this will influence the shape of the deflection profile of the cupula, it is unlikely to give rise to new phenomena.
The model we have presented allows for the cupula to have some pre-tension, $T$, although this has not been observed experimentally. This lack of direct observation is in part because of the great difficulty in accessing this delicate tissue. (Pre-tension is often present in thin biological structures due to residual stresses from development or mechanical interactions with surrounding tissues (Goriely, 2017).) Experiments that determine the presence of pre-tension in the cupula (and its magnitude) could prove useful in understanding the role of parameters like the inverse bendability, $b$.
@@
Another simplification in our model is that that the cupula is uniformly clamped to the crista and canal walls. However, previous results based on this assumption Selva et al. (2009) required a very small Young’s modulus, close to $5\mathrm{~{}Pa}$, to match experimentally observed deformations Selva et al. (2009). This is an extremely low value, perhaps indicating the softest material in the human body, and is unrealistic when compared to other “soft” biological tissues (Goriely, 2017). We suggest that this anomalous stiffness of the cupula may be a result of the clamped boundary conditions on all sides of the cupula, as used here and as usual in the vestibular literature (Rabbitt & Damiano, 1992), may be incorrect; typical anatomical drawings suggest that the cupula is only clamped on a part of its boundary, and is free to move on other regions of the boundary. This would increase the apparent flexibility of the cupula, creating similar system behaviour with without requiring an unusually small Young’s modulus.
Our model may also allow for other phenomena within the vestibular system to be investigated. An interesting avenue using the techniques developed here is the light cupula phenomenon (see Lee et al., 2024, for a review of the concept), as well as related concepts such as the buoyancy hypothesis to explain balance loss after alcohol intake Brandt (1991). These essentially state that when alcohol is consumed, ethanol diffuses faster into the cupula than the surrounding endolymph, changing their density ratio (which under normal functioning is very close to one, so that the cupula is neutrally buoyant). As ethanol is less dense than water, the cupula would then become negatively buoyant, deforming differently than ordinarily and sending incorrect signals to the nervous system. To account for this in our model, we would need to include a buoyancy term so that the cupula can float or sink through the endolymph.
@@
Dominic Vella, https://orcid.org/$0$-0003-1341-8863
Derek Moulton, https://orcid.org/$0$-0003-3597-7973
Appendix A Details for numerical solutions
@@
Figure 12: Numerically obtained deformation of the solid material at $\hat{t}=0.25$ s, normalized by the maximum deformation. Left panel: deformation along the direction perpendicular to the canal centre-line (in the model’s coordinate system this is along $r$). We plot the membrane limit, $b\to 0$ (dashed black line) and the plate limit $b\to\infty$ (dotted line) for comparison. Middle panel: deformation in the direction parallel to the centreline (i.e. along $s$). Note the small change in deformation along $s$, suggesting a plate equation model that only tracks the deformation of the centre of the cupula is appropriate. Right panel: deformation at the centre of the cupula plotted as a function of the Young’s modulus.
In Figure 12 we plot the deformation of the cupula, by showing (left panel) the deformation of the cupula in the direction normal to the flow and (middle panel) the deformation of the cupula in the direction along the flow. In the first case, we observe that the deformation is large in the centre of the structure and zero on the edges, as expected. However, the structure does not seem to be clamped as might have been expected — the edge exhibits a finite gradient. This disagreement might be an effect of the finite thickness of the cupula used in the numerical simulations, or alternatively, might occur because of numerical artifacts and lack of resolution near the edge. For completeness, we have also included the analytical solution of the plate case ($b\to\infty$) with pinned boundary conditions. This solution does not match the numerical solution either, even if it has a finite gradient at the boundary. Following physical intuition and previous work (Rabbitt & Damiano, 1992), we have opted to present the clamped solution in the main text.

### v2

@@ 7.1 Implications and applications
Here, it is worth considering the distinction between the soft cupula ($\kappa\ll 1$) and stiff cupula ($\kappa\gg 1$) regimes in terms of dimensional quantities. Recall that $\kappa$ is defined as $\kappa=E\mathcal{T}a/(R\mu)$, where $\mathcal{T}$ is the timescale for the head motion, $a$ and $R$ are respectively the small and large radii defining the canals, and $E$ is the cupula’s Young’s modulus. We see that the soft cupula regime may be attained for fast movements (small $\mathcal{T}$), highly viscous media ($\mu$ large) and of course soft cupulas in absolute terms (small $E$). The converse holds for the stiff cupula regime. Inserting parameter values for $a$, $R$, $E$ and $\mu$ into the transition value predicted by our model, $\kappa_{c}=32\pi I_{4}\alpha_{0}(\beta)$, we obtain a critical value
$\mathcal{T}=\frac{\kappa_{c}R\mu}{Ea}$ (65)
which may be interpreted as a critical timescale of head motion below which the system responds to angular velocity, and above which the system responds to angular acceleration. Inserting typical values for a human adult, we compute a transition frequency of 0.27 Hz. Interestingly, human experiments with controlled oscillation frequencies have reported a maximum motion sickness when the frequency is around 0.2 Hz (Golding et al., 2001). Our analysis suggests an intriguing possible explanation for this maximal sickness at intermediate frequencies: intermediate frequencies correspond to motion for which the response of the cupular system follows neither the angular acceleration nor velocity. The “neural mismatch” hypothesis predicts that motion sickness is induced in situations where there is a disagreement between visual or vestibular cues and the information anticipated by the nervous system (Benson, 1990). Since the vestibular system may be expected to provide information that matches neither the true acceleration or the true velocity around the transition point we suggest that it may be the transition point between small and large $\kappa$ that causes motion sickness.
@@
A natural question to ask is whether a similar system could be conceived to detect linear acceleration, rather than rotational motion. This might be expected, for instance, in the limiting case where $\epsilon=0$, in which the velocity profile must remain exactly symmetric. In this case, the dimensional cupular deflection can be expressed as
$\hat{\eta}\sim a^{2}R\Omega_{0}/\nu=\epsilon^{2}R^{3}\Omega_{0}/\nu\rightarrow 0,\quad\text{as}\quad\epsilon\rightarrow 0$ (66)
for fixed $R$ (which must remain fixed to maintain physiological realism). Thus, in the limit $\epsilon=0$, the semicircular canals become ineffective at detecting angular motion. Consistent with this prediction of our model, linear motion of the head is indeed detected by a different component of the vestibular system—the otolith organs—which rely on solid inertia (Rabbitt et al., 2004).
Ageing significantly impairs balance and vestibular function, as evidenced by widespread declines in vestibulo‑ocular reflexes, hair cell counts, and neural processing with advancing age (Anson & Jeka, 2016). While the hypothesis that cupular stiffening contributes to vestibular decline is plausible, no direct physiological or medical evidence currently links ageing to such mechanical changes. Instead, clinical data indicate that balance deficits in older adults are more often driven by hormonal or neurochemical modulation and sensory cell loss (Iwasaki & Yamasoba (2014) report 25% fewer hair cells in nonagenarians versus individuals in their 50s). Nonetheless, the idea of altered cupular mechanics remains an interesting direction for future research, and the model we have presented forms an attractive framework for investigating hypotheses and linking with clinical observations.
@@ 7.2 Limitations and future directions
Our model has been based on a number of approximations that have facilitated the analysis that we have presented here. Of these, perhaps the most important is our use of a disk shaped model of the cupula — this is mathematically convenient but is likely to be an over-simplification of the true geometry of the cupula (Selva et al., 2009). Related to the previous point, the semicircular canals (SCCs) are modelled with circular cross-sections, although anatomical observations indicate they are elliptical (Curthoys & Oman, 1987). While the leading-order equations can be readily adapted to more general cross-sectional geometries, this is not straightforward at $\mathcal{O}(\epsilon)$. At the same time, we note that our existing analysis assumes that the cupula’s thickness is constant, but photographic evidence suggests this is likely not the case (Rabbitt et al., 2004). In particular, the cupula seems to be thinner in the centre and thicker towards the edge. Whilst this will influence the shape of the deflection profile of the cupula, it is unlikely to give rise to new phenomena.
@@
Another simplification in our model is that that the cupula is uniformly clamped to the crista and canal walls. However, previous results based on this assumption Selva et al. (2009) required a very small Young’s modulus, close to $5\mathrm{~Pa}$, to match experimentally observed deformations Selva et al. (2009). This is an extremely low value, perhaps indicating the softest material in the human body, and is unrealistic when compared to other “soft” biological tissues (Goriely, 2017). We suggest that this anomalous stiffness of the cupula may be a result of the clamped boundary conditions on all sides of the cupula, as used here and as usual in the vestibular literature (Rabbitt & Damiano, 1992), may be incorrect; typical anatomical drawings suggest that the cupula is only clamped on a part of its boundary, and is free to move on other regions of the boundary. This would increase the apparent flexibility of the cupula, creating similar system behaviour with without requiring an unusually small Young’s modulus. Moreover, materials with similarly low Young’s moduli are often viscoelastic (see, for example, hydrogels in Ahearne et al., 2005). Incorporating a viscoelastic term into our linear model of cupular deformation is in principle straightforward. However, the absence of specific data on the cupula and its material properties would significantly complicate parameter tuning. Additionally, since the system already includes a source of damping through fluid viscosity, disentangling the respective contributions of fluid and solid damping would be non-trivial.
[…]
[Supplementary data]Supplementary material and movies are available at
https://doi.org/10.1017/jfm.2019…
@@
Dominic Vella, https://orcid.org/0000-0003-1341-8863
Derek Moulton, https://orcid.org/0000-0003-3597-7973
Appendix A Angular velocity of the SCC
We consider the motion of a (rigid-body) human standing vertically on a rotating object, such that the person (and hence the SCC) are not positioned on the axis of rotation. For instance, we could have in mind being on a merry-go-round or going around a corner in a vehicle (which are common situations that lead to motion sickness). In this appendix we show that the angular velocity of the SCC is the same as that of the rotating object.
The centre of the human head has position $\mathbf{R}(t)=R\mathbf{e}_{r}$ measured from the centre of the rotating object (see figure 12). The human is rotating with angular velocity $\boldsymbol{\Omega}(t)=\Omega(t)\mathbf{e}_{z}$ about this point, so that $\dot{\boldsymbol{R}}=\boldsymbol{\Omega}\times\mathbf{R}$. Then, $\dot{\boldsymbol{R}}=R\Omega\,\mathbf{e}_{\theta}$. Now, we consider a horizontal semicircular canal, whose central location is at position $\mathbf{R}+\mathbf{r}$, where $\mathbf{r}$ is a position vector from the centre of the head to the canal centre (see figure 12). We can write $\mathbf{r}=r(\cos\varphi\mathbf{e}_{r}+\sin\varphi\mathbf{e}_{\theta})$ for some $\varphi$. If the body is rotating with the moving object, then it is effectively rigid body rotation, so that $\dot{r}=\dot{\varphi}=0$. Since the frame $\{\mathbf{e}_{r},\mathbf{e}_{\theta}\}$ rotates with angular velocity $\Omega(t)\mathbf{e}_{z}$, it follows that
$\dot{\mathbf{r}}=\Omega\mathbf{e}_{z}\times\mathbf{r},$ (67)
meaning that the vector $\mathbf{r}$ undergoes rotation with the same angular velocity as the rotating object in which it sits. Similarly, if we define $\mathbf{x}$ as a vector pointing from the centre of the semicircular canal to a point inside the canal (see figure 12), we have $\dot{\mathbf{x}}=\boldsymbol{\Omega}\times\mathbf{x}$. Writing the position of a part of the semicircular canal as measured from a lab frame as $\mathbf{p}=\mathbf{R}+\mathbf{r}+\mathbf{x}$, we find
$\dot{\mathbf{p}}=\dot{\mathbf{R}}+\dot{\mathbf{r}}+\dot{\mathbf{x}}=\boldsymbol{\Omega}\times\mathbf{p},$ (68)
and hence there is no additional linear velocity.
[…]
Figure 12: Diagram for rotational motion not centered on the SCCs.
Appendix B On the deformation of the cupula
In this section we discuss the details for the numerical simulations from figure 3, and present a method to obtain a polynomial approximation for the deformation of the cupula.
B.1 Details for numerical solutions
@@
Figure 13: Numerically obtained deformation of the solid material at $\hat{t}=0.25$ s, normalized by the maximum deformation. Left panel: deformation along the direction perpendicular to the canal centre-line (in the model’s coordinate system this is along $r$). Middle panel: deformation in the direction parallel to the centreline (i.e. along $z$). Right panel: deformation at the centre of the cupula plotted as a function of the Young’s modulus. All deformations shown have been averaged over the azimuthal direction $\theta$.